OpenStudy (diyadiya):

if in an A.P 3rd term is 18 and 17th term is 30. Sum of its 17 terms is: (I'm not getting the right answer)

OpenStudy (diyadiya):

3rd Term= a+2d 17thTerm=a+16d

OpenStudy (experimentx):

18 = a+2d; 30 =a+16d solve them, find a and d, use in formula of Sn to get value http://www.wolframalpha.com/input/?i=18+%3D+a%2B2d%3B+30+%3Da%2B16d

OpenStudy (diyadiya):

Ok so solution given in my book is wring :)

OpenStudy (callisto):

3rd Term= a+2d =18 -(1) 17thTerm=a+16d =30 -(2) (2)-(1) 14d = 12 d = 12/14 = 6/7 a + 2(6/7) =18 a = 114/7 Sum = (114/7 + 30) x17 /2 = something ugly :S

OpenStudy (experimentx):

i think we should have a calculator applet

OpenStudy (diyadiya):

Yup! Something ugly But the answer given in the book was 612 So its wrong =) Thank you :)

OpenStudy (anonymous):

... isn't that ugly ? LOL

OpenStudy (diyadiya):

Definitely :)